OpenStudy (anonymous):

Find the limit as x approaches 0. 1-cos3x/sin3x.

OpenStudy (anonymous):

use l'hospitals

OpenStudy (anonymous):

what is that?

Parth (parthkohli):

It's a rule in which you differentiate the numerator and denominator to simplify a limit when the fraction is in the form of \(0 \over 0\) or \(\infty\over \infty\)

OpenStudy (anonymous):

how do you use that rule in this problem?

Parth (parthkohli):

You will keep differentiating the numerator and denominator till you stop getting \(0 \over 0\) or \(\infty \over \infty\).

OpenStudy (anonymous):

I haven't learned what differentiating is

OpenStudy (cruffo):

I may be breaking some math laws here but what it you divide both the top and bottom by 3x: \[\large \frac{\frac{1-cos3x}{3x}}{\frac{sin3x}{3x}}\] Then use the special limits for sine and cosine

hartnn (hartnn):

lets do it other way then,  now just put x=0,and evaluate, did u get that? can u solve further?